# Maël Dumas LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Anthony Perez LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Mathis Rocton Algorithms and Complexity Group, TU Wien, Vienna, Austria Ioan Todinca LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France

### v3

@@ 4.4 Bounding the size of the kernel
Proof.
We rely on the linear time computation of the critical clique graph $\mathcal{C}(G)$ of $G$ [38] and the linear time recognition of block graphs. We show that we can enumerate all $1$-SC-branches and $2$-SC-branches in polynomial time. Since an attachment point is by definition a critical clique of $G$, one can detect $1$-SC-branches by removing a critical clique of $\mathcal{C}(G)$ and looking among the remaining connected components those that induce a connected block graph together with $P$ (in $\mathcal{C}(G)$). Considering a maximal set of such components together with $P$ gives a $1$-SC-branch $B$. We proceed similarly to detect clean $2$-SC-branches by removing a pair of critical cliques $P_{1},P_{2}$ of $\mathcal{C}(G)$, and looking among the remaining connected components those that induce a connected block graph (in $\mathcal{C}(G)$) together with $P_{1}$ and $P_{2}$. Such components together with $P_{1}$ and $P_{2}$ gives a clean $2$-SC-branch $B$. Since there is $O(|V(G)|)$ critical cliques, the result follows. ∎
@@
Proof.
Let $(G=(V,E),k)$ be a reduced yes-instance of Strictly Chordal Editing, $F$ a $k$-edition of $G$ and $H=G\triangle F$. We assume that $G$ is connected, the following arguments can be easily adapted if this is not the case by summing over all connected components of $G$. The graph $H$ is strictly chordal, thus the graph $\mathcal{C}(H)$ is a block graph. We first show that $\mathcal{C}(H)$ has $O(k^{3})$ vertices. We say that a critical clique of $\mathcal{C}(H)$ is affected if it contains a vertex affected by $F$. Let $A$ be the set of affected critical cliques of $\mathcal{C}(H)$. Since $|F|\leqslant k$, we have $|A|\leqslant 2k$. Let $T_{A}$ be the minimal connected induced subgraph of $H$ that spans all critical cliques of $A$, and $d(T_{A})$ the set of vertices of degree at least $3$ in $T_{A}$ (see Figure 9). Recall that such a subgraph is unique by Lemma 5.
We first show that $|V(T_{A})|$ contains $O(k^{2})$ critical cliques. By Lemma 5, $|d(T_{A})\setminus A|\leqslant 3\cdot|A|\leqslant 6k$. The connected components of the graph $T_{A}\setminus(A\cup d(T_{A}))$ are paths since every vertex is of degree at most $2$ and by Lemma 5 there are at most $4k$ such paths. Let $R$ be one of these paths, it is composed of unaffected critical cliques, thus there exists a clean $2$-SC-branch $B$ of $G$ that contains $R$. Moreover, the endpoints of $R$ are the attachment points of $B$, hence $B$ have been reduced by Rule 4.5. Thus $R$ is of length at most $k+3$. It remains that $|V(T_{A})|\leqslant 2k+6k+(k+3)\cdot 4k=O(k^{2})$.
@@
We conclude that $|V(\mathcal{C}(H))|=O(k^{3})$. Each unaffected critical clique of $H$ contains $k+1$ vertices by Rule 4.2. Moreover, each affected critical clique of $H$ contains at most $k+1$ unaffected vertices since they are clique modules of $G$ (in particular, they were contained in a critical clique of $G$ which was reduced by Rule 4.2). Therefore, taking into accounts the at most $2k$ affected vertices, we have that $|V(G)|\leqslant(k+1)|V(\mathcal{C}(H))|+2k=O(k^{4})$. Therefore, given a reduced instance $(G,k)$ of Strictly Chordal Editing, if $|V(G)|>g(k)$ for a function $g(k)=O(k^{4})$ (defined by this proof) we can safely conclude that $(G,k)$ is a no-instance and return a trivial no-instance of constant size, thus proving the size of our kernel. To finish the proof, we claim that a reduced instance can be computed in polynomial time. Indeed, Lemma 14 states that it is possible to reduce exhaustively a graph under Rules 4.1 and 4.2 in linear time. Once this is done, it remains to apply exhaustively Rules 4.3, 4.4 and 4.5 which is ensured by Lemma 23. ∎
@@ 5 Kernels for the completion and deletion variants
In this section we present kernels for Strictly Chordal Completion and Strictly Chordal Deletion. It is clear that Rules 4.1 and 4.2 are also safe for these variants of the problem. We can also observe that Rules 4.3, 4.4 and 4.5 are safe for the completion and deletion variants since they are a special case of the edition. However, we can use more efficient rules to manage the $2$-SC-branches and bound their length to a constant number of critical cliques. This is due to only allowing edge addition or edge deletion. In the case of Strictly Chordal Completion, one can observe that for a graph $G$, it is sufficient to consider the clean $2$-SC-branches $B$ of length $3$ which have attachment points that belong to different connected components of $G\setminus V(B^{R})$ since a SC-branch remains connected in any completion of $G$. The next lemma follows from this observation and its proof is contained in the proof of Lemma 21. From this lemma we can devise an improved rule for the completion variant.
Lemma 24.
@@
In the case of Strictly Chordal Deletion, we can observe that if the two attachment points of $B$, a $2$-SC-branch of length at least $3$, are in the same connected component of $G\setminus V(B^{R})$, then either $B$ is disconnected or the attachment points are not in the same connected component of $(G-F)\setminus V(B^{R})$ for any deletion $F$ of $G$. Indeed, if this was not the case, we could find an induced cycle of length at least $4$ since edge additions are not allowed. The next lemma follows from this observation and its proof is contained in the proof of Lemma 21. As previously we can devise an improved rule from this lemma for the deletion variant.
Lemma 25.
@@ 6 Conclusion
We presented polynomial kernels for the edition and deletion variants of block graph edge modification problems and the three variants of strictly chordal edge modification problems. Our conviction is that the approach based on decompositions of the target class, combined with the ability of reducing the size of the bags of the decomposition and of limiting the number of affected bags to $O(k)$ is a promising starting point for edge modification problems, especially into subclasses of chordal graphs. The technique has been employed especially for classes that admit a tree-like decompositions with disjoint bags (e.g., 3-leaf power [3], trivially perfect graphs [22, 20] and $\mathcal{M}$-free graphs [32]), also for other types of tree-like decompositions with non-disjoint bags (e.g., ptolemaic graphs [13]). We generalize it here to strictly chordal graphs, that have a decomposition into disjoint bags as nodes of a block graph.
The difficulty is that, at this stage, each class requires ad-hoc arguments and reduction rules, based on its specific decomposition. An ambitious goal would be to obtain a generic algorithm for edge modification problems into any class of chordal graphs, plus a finite set of obstructions, as conjectured by Bessy and Perez [4] for completion problems. As an intermediate step, we ask whether $4$-leaf power completion, deletion and editing problems admit a polynomial kernel.
On another level, it would be really interesting to obtain polynomial kernelization lower bounds for edge-modification problems. To the best of our knowledge, such bounds are only known for vertex-deletion problems; more precisely, any such problem where the target class is closed under subgraphs does not admit a kernel with $O(k^{2-\epsilon})$ edges unless $coNP\subseteq NP/poly$, see [15].

### v4

@@ 4.4 Bounding the size of the kernel
Proof 4.47.
We rely on the linear time computation of the critical clique graph $\mathcal{C}(G)$ of $G$ (38) and the linear time recognition of block graphs. We show that we can enumerate all $1$-SC-branches and $2$-SC-branches in polynomial time. Since an attachment point is by definition a critical clique of $G$, one can detect $1$-SC-branches by removing a critical clique of $\mathcal{C}(G)$ and looking among the remaining connected components those that induce a connected block graph together with $P$ (in $\mathcal{C}(G)$). Considering a maximal set of such components together with $P$ gives a $1$-SC-branch $B$. We proceed similarly to detect clean $2$-SC-branches by removing a pair of critical cliques $P_{1},P_{2}$ of $\mathcal{C}(G)$, and looking among the remaining connected components those that induce a connected block graph (in $\mathcal{C}(G)$) together with $P_{1}$ and $P_{2}$. Such components together with $P_{1}$ and $P_{2}$ gives a clean $2$-SC-branch $B$. Since there is $O(|V(G)|)$ critical cliques, the result follows.
@@
Proof 4.48.
Let $(G=(V,E),k)$ be a reduced yes-instance of Strictly Chordal Editing, $F$ a $k$-edition of $G$ and $H=G\triangle F$. We assume that $G$ is connected, the following arguments can be easily adapted if this is not the case by summing over all connected components of $G$. The graph $H$ is strictly chordal, thus the graph $\mathcal{C}(H)$ is a block graph. We first show that $\mathcal{C}(H)$ has $O(k^{3})$ vertices. We say that a critical clique of $\mathcal{C}(H)$ is affected if it contains a vertex affected by $F$. Let $A$ be the set of affected critical cliques of $\mathcal{C}(H)$. Since $|F|\leqslant k$, we have $|A|\leqslant 2k$. Let $T_{A}$ be the minimal connected induced subgraph of $H$ that spans all critical cliques of $A$, and $d(T_{A})$ the set of vertices of degree at least $3$ in $T_{A}$ (see Figure 9). Recall that such a subgraph is unique by Lemma 2.10.
We first show that $|V(T_{A})|$ contains $O(k^{2})$ critical cliques. By Lemma 2.10, $|d(T_{A})\setminus A|\leqslant 3\cdot|A|\leqslant 6k$. The connected components of the graph $T_{A}\setminus(A\cup d(T_{A}))$ are paths since every vertex is of degree at most $2$ and by Lemma 2.10 there are at most $4k$ such paths. Let $R$ be one of these paths, it is composed of unaffected critical cliques, thus there exists a clean $2$-SC-branch $B$ of $G$ that contains $R$. Moreover, the endpoints of $R$ are the attachment points of $B$, hence $B$ have been reduced by Rule 4.5. Thus $R$ is of length at most $k+3$. It remains that $|V(T_{A})|\leqslant 2k+6k+(k+3)\cdot 4k=O(k^{2})$.
@@
We conclude that $|V(\mathcal{C}(H))|=O(k^{3})$. Each unaffected critical clique of $H$ contains $k+1$ vertices by Rule 4.2. Moreover, each affected critical clique of $H$ contains at most $k+1$ unaffected vertices since they are clique modules of $G$ (in particular, they were contained in a critical clique of $G$ which was reduced by Rule 4.2). Therefore, taking into accounts the at most $2k$ affected vertices, we have that $|V(G)|\leqslant(k+1)|V(\mathcal{C}(H))|+2k=O(k^{4})$. Therefore, given a reduced instance $(G,k)$ of Strictly Chordal Editing, if $|V(G)|>g(k)$ for a function $g(k)=O(k^{4})$ (defined by this proof) we can safely conclude that $(G,k)$ is a no-instance and return a trivial no-instance of constant size, thus proving the size of our kernel. To finish the proof, we claim that a reduced instance can be computed in polynomial time. Indeed, Lemma 4.28 states that it is possible to reduce exhaustively a graph under Rules 4.1 and 4.2 in linear time. Once this is done, it remains to apply exhaustively Rules 4.3, 4.4 and 4.5 which is ensured by Lemma 4.46.
@@ 5 Kernels for the completion and deletion variants
In this section we present kernels for Strictly Chordal Completion and Strictly Chordal Deletion. It is clear that Rules 4.1 and 4.2 are also safe for these variants of the problem. We can also observe that Rules 4.3, 4.4 and 4.5 are safe for the completion and deletion variants since they are a special case of the edition. However, we can use more efficient rules to manage the $2$-SC-branches and bound their length to a constant number of critical cliques. This is due to only allowing edge addition or edge deletion. In the case of Strictly Chordal Completion, one can observe that for a graph $G$, it is sufficient to consider the clean $2$-SC-branches $B$ of length $3$ which have attachment points that belong to different connected components of $G\setminus V(B^{R})$ since a SC-branch remains connected in any completion of $G$. The next lemma follows from this observation and its proof is contained in the proof of Lemma 4.42. From this lemma we can devise an improved rule for the completion variant.
Lemma 5.49.
@@
In the case of Strictly Chordal Deletion, we can observe that if the two attachment points of $B$, a $2$-SC-branch of length at least $3$, are in the same connected component of $G\setminus V(B^{R})$, then either $B$ is disconnected or the attachment points are not in the same connected component of $(G-F)\setminus V(B^{R})$ for any deletion $F$ of $G$. Indeed, if this was not the case, we could find an induced cycle of length at least $4$ since edge additions are not allowed. The next lemma follows from this observation and its proof is contained in the proof of Lemma 4.42. As previously we can devise an improved rule from this lemma for the deletion variant.
Lemma 5.50.
@@ 6 Conclusion
We presented polynomial kernels for the edition and deletion variants of block graph edge modification problems and the three variants of strictly chordal edge modification problems. Our conviction is that the approach based on decompositions of the target class, combined with the ability of reducing the size of the bags of the decomposition and of limiting the number of affected bags to $O(k)$ is a promising starting point for edge modification problems, especially into subclasses of chordal graphs. The technique has been employed especially for classes that admit a tree-like decompositions with disjoint bags (e.g., 3-leaf power (3), trivially perfect graphs (22, 20) and $\mathcal{M}$-free graphs (32)), also for other types of tree-like decompositions with non-disjoint bags (e.g., ptolemaic graphs (13)). We generalize it here to strictly chordal graphs, that have a decomposition into disjoint bags as nodes of a block graph.
The difficulty is that, at this stage, each class requires ad-hoc arguments and reduction rules, based on its specific decomposition. An ambitious goal would be to obtain a generic algorithm for edge modification problems into any class of chordal graphs, plus a finite set of obstructions, as conjectured by Bessy and Perez (4) for completion problems. As an intermediate step, we ask whether $4$-leaf power completion, deletion and editing problems admit a polynomial kernel.
On another level, it would be really interesting to obtain polynomial kernelization lower bounds for edge-modification problems. To the best of our knowledge, such bounds are only known for vertex-deletion problems; more precisely, any such problem where the target class is closed under subgraphs does not admit a kernel with $O(k^{2-\epsilon})$ edges unless $coNP\subseteq NP/poly$, see (15).